1. Consider the following economy of Syldavia (a small open economy)
Y=C+I+G+NX , NX = S-I
Y=8000
G=750
T=750
C=1000+0.75(Y-T)
I=1000-100r
NX=500-500e
r=r*=5

d. [ 5 points] Suppose the world interest rate drop from r=5 to 2percent (assume government
G=750). Find the national saving, investment, trade balance, capital outflow and equilibrium
exchange rate.
Answer #1

(I)

National saving = (Y - C) + (T - G)

= 8000 - [1000 + 0.75 x (8000 - 1000)] + (750 - 750)

= 8000 - [1000 + (0.75 x 7000)] + 0

= 8000 - (1000 + 5250)

= 8000 - 6250

= 1750

(II)

I = 1000 - (100 x 5)

= 1000 - 500

= 500

(III)

Trade balance (X - M = NX) = (S - I) where S: National saving

= 1750 - 500

= 1250

(IV)

Capital outflow = Trade balance = NX

= 1250

(V)

NX = 500 - 500e

1250 = 500 - 500e

500e = 750

e = 1.5